only choose mcq Four resistors, A: 25 ohm, B: 50 ohm, C: 10 ohm and D:...

  1. only choose mcq

  2. Four resistors, A: 25 ohm, B: 50 ohm, C: 10 ohm and D: 80 ohm are connected in series. A battery of 10 V is applied across the combination. If current is passed through the combination for one hour, heat generated is minimum for which resistor?

    A.

    C

    B.

    A

    C.

    D

    D.

    B
Answer #1

Answer is Resistor C

equivalent resistance is given by

R=25+50+10+80

R=165 A

Current flowing in the circuit is

I=V/R =10/165 =0.060606 A

Heat generated in each resistor is

QA =I2RAt =0.0606062*25*(1*60*60) =330.58 J

QB =I2RBt =0.0606062*50*(1*60*60) =667.16 J

QC =I2RCt =0.0606062*10*(1*60*60) =132.23 J

QD=I2RDt =0.0606062*80*(1*60*60) =1057.85 J

Therefore answer is A. C